OpenStudy (anonymous):

the ratio if the base edges of two similar pyramids is 3:4 the volume of the larger pyramid is 320 in^3. what is the volume of the smaller pyramid???

OpenStudy (anonymous):

ratio of voljumes = 3^3 : 4^3 or 27:64 small/large = 27 / 64 vol of smaller = 27 * 320 / 64 =135